find the no which when divided by 17 ,16,15 and the remainder is 3.10,0 respectively

Grade:10

2 Answers

Aman Bansal
12 years ago

Daer Sathyaram,

use the chinese remainder theorem for this.

it is helpful in these types of questions.

we want to solve.

x = 0 (mod 15).

x = 3 (mod 17).

and x = 10 (mod 16).

a = b(modn) means that

(a-b) is divisible by n.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Sathya
12 years ago

Sir,

 

I know wht u posted sir....

 

But i couldnt proceed...

 

Pls help me